A cat chases a mouse across a 1.3 m high
table. The mouse steps out of the way, and
the cat slides off the table and strikes the floor
2.1 m from the edge of the table.
The acceleration of gravity is 9.81 m/s
2. What was the cat’s speed when it slid off
the table?
Answer in units of m/s. Don't round answer.
1
Expert's answer
2019-10-10T09:42:56-0400

The time of flight:


"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(1.3)}{9.81}}=0.515\ s"

The cat’s speed when it slid off the table is


"v=\frac{x}{t}=\frac{2.1}{0.515}=4.1\frac{m}{s}"